# Gravitational Waves from Type-I Strings in a Neutrino Mass Model

### v1

@@ Abstract
In this work, we propose a novel realization of type-I cosmic strings arising from the spontaneous breaking of an extended gauge symmetry $SU(2)_{R}\times U(1)_{B-L}$ in the context of a low-scale split seesaw mechanism for neutrino mass generation. We demonstrate that the split seesaw framework, which explains the smallness of neutrino masses, naturally motivates a small scalar self-coupling $\lambda$. This intrinsically links the neutrino mass generation mechanism to the formation of type-I cosmic strings, where the gauge coupling dominates over the scalar self-coupling ($\beta\equiv\lambda/2g^{2}<1$). We explore the cosmological implications of these strings, including their gravitational wave signatures that are testable in current and future experiments. Our findings establish a compelling and testable connection between neutrino mass generation and cosmic string phenomenology in an underexplored region of parameter space.
@@ I Introduction
offer a compelling framework to address these questions. They naturally incorporate right-handed (RH) neutrinos as gauge singlets under the Standard Model, but charged under $SU(2)_{R}\times U(1)_{B-L}$, and thus enable the implementation of the seesaw mechanism for neutrino mass generation [1, 2]. Moreover, they provide a natural origin for parity violation in weak interactions, embedding it within a spontaneously broken symmetry [3].
@@
Among the various implementations of the seesaw mechanism, the split seesaw scenario [4] provides an elegant explanation for the tiny active neutrino masses through geometric suppression in higher-dimensional theories. In this framework, RH neutrinos are localized at different positions in a compactified extra dimension. The overlap of their wavefunctions with the Standard Model brane results in exponentially suppressed Yukawa couplings, without requiring ultra-small fundamental parameters. This separation not only generates a realistic neutrino mass spectrum but also enables one of the RH neutrinos to be long-lived and a viable dark matter candidate [5, 6].
@@
A critical prediction of LR symmetric theories is the spontaneous breaking of the gauge symmetry $SU(2)_{R}\times U(1)_{B-L}\rightarrow U(1)_{Y}$, which is typically achieved by the vacuum expectation value (VEV) of a RH triplet scalar $\Delta_{R}$. This symmetry breaking can lead to the formation of topological defects, specifically cosmic strings (CSs), if the vacuum manifold contains non-trivial first homotopy groups [7, 8]. In our model, this breaking gives rise to type-I CSs, characterized by a scalar self-coupling $\lambda$ much smaller than the gauge coupling squared $g^{2}$, i.e., $\beta\equiv\lambda/g^{2}\ll 1$ [9]. The formation of type-I CSs from supersymmetric flat directions has been explored in [10]. Such strings are attractive in nature and can influence early-universe dynamics and gravitational wave (GW) backgrounds.
@@
Furthermore, the large VEV, $\nu_{R}$ associated with $\Delta_{R}$ not only sets the symmetry-breaking scale but also giving mass to the heavy $U(1)_{B-L}$ gauge boson $Z^{{}^{\prime}}$. In the split seesaw context, we show that the scale $\nu_{R}\simeq(10^{5}-3\times 10^{13})$ GeV that is consistent with both GW observations and evades the collider bounds [11]. The heavy mass of the gauge boson suppresses its emission from the string network, favoring decay channels such as GWs or sterile neutrino emission. This suppression is essential in our model, where we explore the cosmological signatures of such strings without conflicting with astrophysical constraints.
@@
In this work, we analyze the formation and evolution of type-I CSs in the framework of the split seesaw mechanism embedded in an LR symmetric model. We discuss the implications for neutrino mass generation and CS phenomenology, and identify benchmark parameter sets consistent with current theoretical and observational bounds. Our results demonstrate the compatibility of this unified framework with high-energy particle physics and early universe cosmology. This regime is often overlooked in the literature, as most cosmic string studies focus on type-II strings, where parallel strings tend to repel at large distances while antiparallel strings attract [7].
@@
The unique synergy of this framework is threefold: Firstly, the split seesaw provides a natural explanation for small neutrino masses and motivates the small quartic coupling necessary for type-I strings. Secondly, the left-right symmetry-breaking scale $\nu_{R}$ is simultaneously the source of heavy neutrino masses and the string tension. Thirdly, this connects the scale of neutrino mass generation directly to an observable cosmological signal-GWs from CSs—offering a novel, multi-messenger probe of neutrino physics. The leptogenesis in the LR model has been discussed in Ref. [12] and some analysis of GW production in the LR model has been discussed in Ref. [13, 14, 15, 16].
@@
The paper is organized as follows: in Sec. II, we discuss the LR symmetric model and the formation of CSs as a result of symmetry breaking, we define the Lagrangian and scalar potential in Sec. II.1, in Sec. II.2 we discuss the neutrino mass generation in the split seesaw framework. We briefly discuss type-1 CSs and their phenomenological implications in Sec. III and lastly in Sec. IV we summarize our conclusions.
@@ II.1 Lagrangian and Scalar Potential
$\mathcal{L}\supset\text{Tr}[(D_{\mu}\Delta_{R})^{\dagger}(D^{\mu}\Delta_{R})]-V(\Delta_{R}),$ (6)
@@
$D_{\mu}\Delta_{R}=\partial_{\mu}\Delta_{R}-ig_{R}[W_{R\mu},\Delta_{R}]-i2g_{B-L}B_{\mu}\Delta_{R},$ (7)
where $\partial_{\mu}$ is the ordinary derivative, $g_{R},\,W_{R_{\mu}}$ is the associated gauge coupling and bosons with $SU(2)_{R}$ respectively, $g_{B-L},\,B_{\mu}$ is the associated gauge coupling and bosons of $U(1)_{B-L}$. The scalar potential for $\Delta_{R}$ is,
@@
Here $\mu$ is a positive mass-squared parameter that triggers spontaneous symmetry breaking when the potential is minimized, and $\lambda$ is the quartic self-coupling constant. This potential leads to a non-zero VEV for $\Delta_{R}$, breaking $SU(2)_{R}\times U(1)_{B-L}\rightarrow U(1)_{Y}$ and producing Majorana masses for RH neutrinos via the seesaw mechanism [19, 20, 21]. Assuming a vacuum configuration:
@@
$\mu\simeq 2\pi B(\beta)\nu_{R}^{2}.$ (10)
Here $\nu_{R}$ is the VEV of the scalar field $\Delta_{R}$, $\beta\equiv m_{\Delta_{R}}^{2}/m_{W_{R}}^{2}$ is the ratio between the mass square of the scalar and gauge field. The function $B(\beta)$ encodes the dependence on the scalar-to-gauge mass ratio. While its asymptotic behavior is well-known ($B(\beta)\rightarrow 1$ for $\beta\ll 1$ and $B(\beta)\sim ln(\beta)$ for $\beta\gg 1$) [22, 23], we adopt a phenomenological interpolation formula that accurately approximates numerical solutions across the full range of $\beta$ [24, 25]:
$B(\beta)\sim\dfrac{1}{1+2.5\left(\beta^{-1}-1\right)^{1.2}}.$ (11)
This fit is chosen for its simplicity and accuracy in representing the transition between the type-I ($\beta<1$) and type-II ($\beta>1$) regimes, which is crucial for our analysis.
@@ II.2 Mass Generation in Split Seesaw
In split seesaw models, the smallness of the neutrino Yukawa coupling arises from localization in extra dimensions [26, 4]. In the split seesaw framework, we assume the existence of a single extra spatial dimension $y\in[0,L]$. The Standard Model fields are localized on a 4D brane at $y=0$, while the RH neutrinos $N_{i}$ propagate in the bulk. The 5D wavefunction of the $i$-th RH neutrino has the form:
$\Psi_{N_{i}}(x,y)=\psi_{N_{i}}(x)f_{i}(y),\quad\text{with}\quad f_{i}(y)=\sqrt{2M_{i}}e^{-M_{i}y},$ (12)
where $M_{i}$ is the 5D bulk mass term, assumed positive. The 5D Yukawa coupling to the Higgs field $\Phi$ localized at $y=0$ generates a suppressed effective 4D Yukawa coupling:
$Y^{\text{eff}}_{i\alpha}=Y^{(5D)}_{i\alpha}\cdot f_{i}(0)=Y^{(5D)}_{i\alpha}\cdot\sqrt{2M_{i}}e^{-M_{i}L}.$ (13)
The Majorana mass arises from a coupling of $\Psi_{N_{i}}$ to the $\Delta_{R}$ field on the brane at $y=L$:
@@
The Dirac mass term is:
$m_{D}=Y^{\text{eff}}v_{\text{EW}}\sim e^{-M_{i}L}v_{\text{EW}}.$ (15)
The standard type-I seesaw formula then gives:
@@
naturally yielding small active neutrino masses without requiring tiny dimensionless couplings.
@@
The exponential suppression $e^{-M_{i}L}$ requires $L>1/M_{i}$ to be significant. However, to ensure the validity of the 44D effective field theory (EFT) at the temperature of the $SU(2)_{R}\times U(1)_{B-L}$ phase transition ($T\sim\nu_{R}$), we must impose that the compactification scale, $M_{c}\sim 1/L$ is significantly higher than $\nu_{R}$. This avoids Kaluza-Klein mode excitations and higher-dimensional operators disrupting the dynamics of CS formation. Consequently, we require
$\nu_{R}\ll M_{c}<M_{i}.$ (17)
@@
This hierarchy ensures that the 44D description of the phase transition and the resulting topological defects remain robust. The small Yukawa couplings are generated by the large bulk mass terms $M_{i}$, not an excessively large extra dimension that would compromise the EFT. In Fig. 1, we have shown the neutrino mass evolution in the split seesaw mechanism versus the symmetry-breaking scale. The colored shaded region indicates the NANOGrav [27] excluded regime as illustrated in Fig. 2 via frequency bins. The inside colored curves indicate the different values of the exponential suppression factor. The colored stars correspond to the atmospheric scale mass, $M_{\nu}\simeq 0.05$ eV for each exponential suppression factor. The colliders constrain the lower value $\nu_{R}\gtrsim 10^{5}$ GeV [11].
@@
Figure 1: Neutrino mass evolution versus symmetry-breaking scale has been presented. The colored shaded region indicates the NANOGrav [27] excluded regime as shown in Fig. 2. The colored curves represent the different values of the exponential suppression factor. The colored stars correspond to the atmospheric scale mass, $M_{\nu}\simeq 0.05$ eV for each exponential suppression factor. The colliders constrain the lower value $\nu_{R}\gtrsim 10^{5}$ GeV [11].

### v2

@@ Abstract
In this work, we propose a novel realization of type-I cosmic strings arising from the spontaneous breaking of an extended gauge symmetry $SU(2)_{R}\times U(1)_{B-L}$ in the context of a low-scale split seesaw mechanism for neutrino mass generation. We demonstrate that the split seesaw framework, which explains the smallness of neutrino masses, naturally motivates a small scalar self-coupling $\lambda$. This intrinsically links the neutrino mass generation mechanism to the formation of type-I cosmic strings, where the gauge coupling dominates over the scalar self-coupling ($\beta\equiv\lambda/(2g^{2})<1$). We explore the cosmological implications of these strings, including their gravitational wave signatures that are testable in current and future experiments. Our findings establish a compelling and testable connection between neutrino mass generation and cosmic string phenomenology in an underexplored region of parameter space.
@@ I Introduction
offer a compelling framework to address these questions. They naturally incorporate right-handed (RH) neutrinos as gauge singlets under the Standard Model, but charged under $SU(2)_{R}\times U(1)_{B-L}$, and thus enable the implementation of the seesaw mechanism for neutrino mass generation [31, 30]. Moreover, they provide a natural origin for parity violation in weak interactions, embedding it within a spontaneously broken symmetry [37]. The leptogenesis in the LR model has been discussed in Ref. [22] and some analysis of GW production in the LR model has been discussed in Ref. [26, 5, 6, 3]. The study of leptogenesis and GWs from $U(1)_{B-L}$ breaking has been discussed in [13], type-1 CSs in MSSM connecting electron neutrino asymmetry hinted by the recent EMPRESS data have recently been explored in [8].
@@
Among the various implementations of the seesaw mechanism, the split seesaw scenario [25, 2] provides an elegant explanation for the tiny active neutrino masses through geometric suppression in higher-dimensional theories. In this framework, RH neutrinos are localized at different positions in a compactified extra dimension. The overlap of their wavefunctions with the Standard Model brane results in exponentially suppressed Yukawa couplings, without requiring ultra-small fundamental parameters. This separation not only generates a realistic neutrino mass spectrum but also enables one of the RH neutrinos to be long-lived and a viable dark matter candidate [7, 20].
@@
A critical prediction of LR symmetric theories is the spontaneous breaking of the gauge symmetry $SU(2)_{R}\times U(1)_{B-L}\rightarrow U(1)_{Y}$, which is typically achieved by the vacuum expectation value (VEV) of a RH triplet scalar $\Delta_{R}$. This symmetry breaking can lead to the formation of topological defects, specifically cosmic strings (CSs), if the vacuum manifold contains non-trivial first homotopy groups [41, 24]. In our model, this breaking gives rise to type-I CSs, characterized by a scalar self-coupling $\lambda$ much smaller than the gauge coupling squared $g^{2}$, i.e., $\beta\equiv\lambda/(2g^{2})\ll 1$ [18]. The formation of type-I CSs from supersymmetric flat directions has been explored in [15]. Such strings are attractive in nature and can influence early-universe dynamics and gravitational wave (GW) backgrounds.
@@
Furthermore, the large VEV, $v_{R}$ associated with $\Delta_{R}$ not only sets the symmetry-breaking scale but also giving mass to the heavy $U(1)_{B-L}$ gauge boson $Z^{{}^{\prime}}$. In the split seesaw context, we show that the scale $v_{R}\simeq(10^{5}-10^{15})$ GeV that is consistent with both GW observations and evades the collider bounds [33]. The heavy mass of the gauge boson suppresses its emission from the string network, favoring decay channels such as GWs or sterile neutrino emission. This suppression is essential in our model, where we explore the cosmological signatures of such strings without conflicting with astrophysical constraints.
@@
In this work, we analyze the formation and evolution of type-I CSs in the framework of the split seesaw mechanism embedded in an LR symmetric model. We discuss the implications for neutrino mass generation and CS phenomenology and identify benchmark parameter sets consistent with current theoretical and observational bounds. Our results demonstrate the compatibility of this unified framework with high-energy particle physics and early universe cosmology. This regime is often overlooked in the literature, as most cosmic string studies focus on type-II strings, where parallel strings tend to repel at large distances while antiparallel strings attract [41].
@@
The unique synergy of this framework is threefold: Firstly, the split seesaw provides a natural explanation for small neutrino masses and motivates the small quartic coupling necessary for type-I strings. Secondly, the left-right symmetry-breaking scale $v_{R}$ is simultaneously the source of heavy neutrino masses and the string tension. Thirdly, this connects the scale of neutrino mass generation directly to an observable cosmological signal-GWs from CSs—since we have shown that the mass suppression factor is constrained by the choice of $\beta$.
@@
The paper is organized as follows: in Sec. II, we discuss the LR symmetric model and the formation of CSs as a result of symmetry breaking; we define the Lagrangian and scalar potential in Sec. II.1, in Sec. II.2 we discuss the neutrino mass generation in the split seesaw framework. We briefly discuss type-1 CSs and their phenomenological implications in Sec. III and lastly, in Sec. IV we summarize our conclusions.
@@ II.1 Lagrangian and Scalar Potential
$\mathcal{L}\supset\text{Tr}[(D_{A}\Delta_{R})^{\dagger}(D^{A}\Delta_{R})]-V(\Delta_{R}),$ (6)
@@
$D_{A}\Delta_{R}=\partial_{A}\Delta_{R}-ig\,[W_{RA},\Delta_{R}]-i2g_{B-L}B_{A}\Delta_{R},$ (7)
where $\partial_{A}$ is the ordinary derivative, $g,\,W_{RA}$ is the associated gauge coupling and bosons with $SU(2)_{R}$ respectively, $g_{B-L},\,B_{A}$ is the associated gauge coupling and bosons of $U(1)_{B-L}$. The scalar potential for $\Delta_{R}$ is,
@@
Here $\mu$ is a positive mass-squared parameter that triggers spontaneous symmetry breaking when the potential is minimized and $\lambda$ is the quartic self-coupling constant. A complete treatment of the scalar potential in the LRSM requires including cross-couplings between $\Phi$, $\Delta_{L}$, and $\Delta_{R}$ to ensure the desired vacuum alignment. A general renormalizable scalar potential in the minimal LRSM can be written as [11]
[…]
$+\lambda_{2}\left(\text{Tr}[\tilde{\Phi}\Phi^{\dagger}]^{2}+\text{Tr}[\tilde{\Phi}^{\dagger}\Phi]^{2}\right)+\lambda_{3}\text{Tr}[\tilde{\Phi}\Phi^{\dagger}]\text{Tr}[\tilde{\Phi}^{\dagger}\Phi]+\lambda_{4}\text{Tr}[\Phi^{\dagger}\Phi]\left(\text{Tr}[\tilde{\Phi}\Phi^{\dagger}]+\text{Tr}[\tilde{\Phi}^{\dagger}\Phi]\right)$
$+\rho_{1}\left(\text{Tr}[\Delta_{L}\Delta_{L}^{\dagger}]^{2}+\text{Tr}[\Delta_{R}\Delta_{R}^{\dagger}]^{2}\right)+\rho_{2}\left(\text{Tr}[\Delta_{L}\Delta_{L}]\text{Tr}[\Delta_{L}^{\dagger}\Delta_{L}^{\dagger}]+\text{Tr}[\Delta_{R}\Delta_{R}]\text{Tr}[\Delta_{R}^{\dagger}\Delta_{R}^{\dagger}]\right)$
$+\rho_{3}\text{Tr}[\Delta_{L}\Delta_{L}^{\dagger}]\text{Tr}[\Delta_{R}\Delta_{R}^{\dagger}]+\rho_{4}\left(\text{Tr}[\Delta_{L}\Delta_{L}]\text{Tr}[\Delta_{R}^{\dagger}\Delta_{R}^{\dagger}]+\text{Tr}[\Delta_{L}^{\dagger}\Delta_{L}^{\dagger}]\text{Tr}[\Delta_{R}\Delta_{R}]\right)$
$+\alpha_{1}\text{Tr}[\Phi^{\dagger}\Phi]\left(\text{Tr}[\Delta_{L}\Delta_{L}^{\dagger}]+\text{Tr}[\Delta_{R}\Delta_{R}^{\dagger}]\right)+\alpha_{3}\left(\text{Tr}[\Phi\Phi^{\dagger}\Delta_{L}\Delta_{L}^{\dagger}]+\text{Tr}[\Phi^{\dagger}\Phi\Delta_{R}\Delta_{R}^{\dagger}]\right)$
$+\alpha_{2}\left(\text{Tr}[\Delta_{L}\Delta_{L}^{\dagger}]\text{Tr}[\tilde{\Phi}\Phi^{\dagger}]+\text{Tr}[\Delta_{R}\Delta_{R}^{\dagger}]\text{Tr}[\tilde{\Phi}^{\dagger}\Phi]+{\rm h.c.}\right)$
$+\beta_{1}\left(\text{Tr}[\Phi\Delta_{R}\Phi^{\dagger}\Delta_{L}^{\dagger}]+\text{Tr}[\Phi^{\dagger}\Delta_{L}\Phi\Delta_{R}^{\dagger}]\right)+\beta_{2}\left(\text{Tr}[\tilde{\Phi}\Delta_{R}\Phi^{\dagger}\Delta_{L}^{\dagger}]+\text{Tr}[\tilde{\Phi}^{\dagger}\Delta_{L}\Phi\Delta_{R}^{\dagger}]\right)$
$+\beta_{3}\left(\text{Tr}[\Phi\Delta_{R}\text{$\tilde{\Phi}^{\dagger}\Delta_{L}^{\dagger}$}]+\text{Tr}[\Phi^{\dagger}\Delta_{L}\text{$\tilde{\Phi}\Delta_{R}^{\dagger}$}]\right)\,$,
where $\tilde{\Phi}=\sigma_{2}\Phi^{*}\sigma_{2}$ (with $\sigma_{2}$ the second Pauli matrix). The vacuum alignment relevant for our analysis is
$\langle\Phi\rangle=\begin{pmatrix}\kappa_{1}&0\\ 0&\kappa_{2}\end{pmatrix},\quad\langle\Delta_{L}\rangle=0,\quad\langle\Delta_{R}\rangle=\dfrac{1}{\sqrt{2}}\begin{pmatrix}0&0\\ v_{R}&0\end{pmatrix}$, (10)
with $v_{R}\gg\kappa_{1}$, $\kappa_{2}\sim v_{\text{EW}}$. This structure arises in the type‑I seesaw limit of the LRSM, where the cross‑couplings between $\Phi$, $\Delta_{L}$ and $\Delta_{R}$ in the full scalar potential can be chosen small enough that the global minimum satisfies $\langle\Delta_{L}\rangle\approx 0$, $\langle\Phi\rangle\simeq v_{\text{EW}}\ll\langle\Delta_{R}\rangle\simeq v_{R}$ while maintaining vacuum stability [29]. Any induced $\langle\Delta_{L}\rangle$ is then suppressed as $v_{L}\sim v_{\text{EW}}^{2}/v_{R}$ and is negligible for neutrino masses. Crucially, the parameter $\beta=\lambda/(2g^{2})$ depends only on the self‑coupling $\lambda\equiv\rho_{1}$ of $\Delta_{R}$ and the gauge coupling $g$. The LRSM features a two-stage symmetry breaking:
$SU(2)_{L}\times SU(2)_{R}\times U(1)_{B-L}$ $\xrightarrow[]{\langle\Delta_{R}\rangle}SU(2)_{L}\times U(1)_{Y}$ (11)
$\xrightarrow[]{\langle\Phi\rangle}U(1)_{\text{EM}}.$
For the purpose of studying CS formation from the primary symmetry breaking $SU(2)_{R}\times U(1)_{B-L}\to U(1)_{Y}$, we work in the limit where $v_{R}\gg v_{\text{EW}}$. This hierarchy is phenomenologically required and theoretically well-motivated in the standard seesaw mechanism. However, in our framework, neutrino masses are generated via the split seesaw mechanism and to be in the observable regime as shown in Fig. 1, our analysis considers $v_{R}$ in the range $10^{5}-10^{15}$ GeV, still maintains $v_{R}\gg v_{\text{EW}}$. The influence of $\Phi$ and $\Delta_{L}$ on these specific string characteristics is subleading, as their VEVs are much smaller than $v_{R}$. The first breaking in Eq. 11 produces Majorana masses for the RH neutrinos, a key element of the seesaw framework [28, 36, 29].
Assuming a vacuum configuration:
@@
$\mu_{cs}\simeq 2\pi B(\beta)v_{R}^{2}.$ (13)
Here $v_{R}$ is the VEV of the scalar field $\Delta_{R}$, $\beta\equiv m_{\Delta_{R}}^{2}/m_{W_{R}}^{2}$ is the ratio between the mass square of the scalar and gauge field. The function $B(\beta)$ encodes the dependence on the scalar-to-gauge mass ratio. While its asymptotic behavior is well-known ($B(\beta)\sim(ln(1/\beta))^{-1}$ for $\beta\ll 1$, $B(\beta)\rightarrow 1$ for $\beta=1$ and $B(\beta)\sim ln(\beta)$ for $\beta\gg 1$) [43, 39], we adopt a phenomenological interpolation formula that approximates numerical solutions across $\beta\leq 1$:
$B(\beta)\simeq(1+|Log\left(\beta\right)|)^{-1}.$ (14)
This fit is chosen for its simplicity and accuracy in representing the region of our interest ($\beta\leq 1$). This fit reproduces the asymptotic behavior for $\beta\ll 1$ given in [43, 39] to within $5\%$. We have shown the impact of $\beta$ on the CSs in Fig. 2.
@@ II.2 Mass Generation in Split Seesaw
In the minimal LRSM, the neutrino mass matrix arises from the Yukawa interactions and the seesaw mechanism. The relevant Yukawa Lagrangian terms are:
$\mathcal{L}_{Y}\supset-Y_{\ell}\overline{L}_{L}\Phi L_{R}-Y_{D}\overline{L}_{L}\tilde{\Phi}L_{R}-\frac{1}{2}Y_{R}\overline{L_{R}^{c}}\Delta_{R}L_{R}+\text{h.c.}$ (15)
After symmetry breaking the $6\times 6$ neutrino mass matrix in the basis $(v_{L},v_{R})$ can be written as
$M_{\nu}=\begin{pmatrix}m_{L}&m_{D}\\ m_{D}^{T}&M_{R}\end{pmatrix}$, (16)
with $m_{L}=Y_{L}v_{L},m_{D}=Y_{D}v_{\text{EW}}$ and $M_{R}=Y_{R}v_{R}$. Diagonalizing this matrix gives
$M_{\nu}=m_{L}-m_{D}\,M_{R}^{-1}\,m_{D}^{T}.$ (17)
In the minimal LRSM and also in our case as discussed in Sec. II.1, $v_{L}$ is negligible and therefore Eq. 17 reduces to the standard type-I seesaw formula
$M_{\nu}=-m_{D}\,M_{R}^{-1}\,m_{D}^{T}.$ (18)
However, in the full LRSM, the LH triplet $\Delta_{L}$ may also acquire a small induced VEV, $v_{L}$, giving an additional contribution $m_{L}=Y_{L}v_{L}$ to the neutrino mass matrix in Eq. 17. The split seesaw mechanism —where the smallness of the neutrino Yukawa coupling arises from localization in extra dimensions [25, 2] —can be consistently formulated including this term. We assume a fifth dimension $y\in[0,L]$ with the Standard Model fields (including $L_{L},\Phi,\Delta_{L}$) are localized on a 4D brane at $y=0$. The RH neutrinos $N_{i}$ and the scalar triplet $\Delta_{R}$ propagate in the Bulk. Their zero mode wavefunctions are chosen to be localized near the distant brane at $y=L$ in order to suppress couplings to the SM brane. For the leptonic sector, the relevant Yukawa interactions and the 55D kinetic term for the RH neutrinos are
[…]
$\left.\left.-\dfrac{1}{2}Y_{L}\bar{L}_{L}^{c}\Delta_{L}L_{L}+\text{h.c}\right)+\left(\bar{N}i\Gamma^{A}D_{A}N-M_{i}\bar{N}N+\text{h.c}\right)\right$.
$\left.+\left(Y_{R}\bar{N^{c}}\Delta_{R}N+\text{h.c}\right)\right],$
where $N$ is the 55D fermion containing the RH neutrino zero mode, $M_{i}$ is its bulk mass assumed to be positive, $Y_{R}$ is the 55D Yukawa coupling, $D_{A}=\partial_{A}-igW^{a}_{RA}T^{a}-i\,2\,g_{B-L}B_{A}$ ($A=0,1,2,3,4$ and $T^{a}$ are the generators of $SU(2)_{R}$) is the 55D covariant derivative and $\Gamma^{A}=(\gamma^{\mu},i\gamma^{5})$ are the 55D gamma matrices. The 55D wavefunction of the $i$-th RH neutrino, localized near the brane at $y=L$, takes the form:
$\Psi_{N_{i}}(x,y)=\psi_{N_{i}}(x)f_{i}(y),\quad\text{with}\quad f_{i}(y)\simeq\sqrt{2M_{i}}\,e^{-M_{i}|y-L|}.$ (20)
Similarly, the bulk scalar field $\Delta_{R}$ has a wavefunction profile
$\Delta_{R}(x,y)=\Phi_{\Delta_{R}}(x)f_{\Delta}(y),\quad\quad f_{\Delta}(y)\simeq\sqrt{2M_{\Delta}}\,e^{-M_{\Delta}|y-L|},$ (21)
where $M_{\Delta}>0$ localizes $f_{\Delta}(y)$ away from $y=0$. The 5D Yukawa coupling to the Higgs field $\Phi$ localized at $y=0$ generates a suppressed effective 4D Yukawa coupling:
$Y^{\text{eff}}_{i\alpha}=Y^{(5D)}_{i\alpha}\cdot f_{i}(0)\propto Y^{(5D)}_{i\alpha}\cdot e^{-M_{i}L}.$ (22)
After dimensional reduction, the effective 44D Dirac and Majorana masses are
$m_{D}$ $=Y^{\text{eff}}_{D}v_{\text{EW}}\sim e^{-M_{i}L}v_{\text{EW}}.$ (23)
The Majorana mass arises from a coupling of $\Psi_{N_{i}}$ to the $\Delta_{R}$ field in the bulk. For the Yukawa interaction, the wavefunction overlap integral is dominated by the region near $y=L$, giving
@@
where $v_{R}=\langle\Phi_{\Delta_{R}}\rangle$ is the 44D VEV. We assume that $\Delta_{L}$ is localized at $y=0$, therefore, $m_{L}=Y_{L}v_{L}$. The neutrino mass from Eq. 17 is
$M_{\nu}\simeq Y_{L}v_{L}+e^{-2M_{i}L}\dfrac{v_{\text{EW}}^{2}}{v_{R}}.$ (25)
For phenomenological consistency of the LRSM (electroweak precision, neutrino masses, and collider bounds) requires the hierarchy $v_{R}\gg\kappa_{1,2}\gg v_{L}$ [16]. In our split‑seesaw realization, $v_{R}$ is the high‑scale VEV of $\Delta_{L}$ localized on a distant brane, while $v_{L}$ is suppressed both by the LRSM scalar potential (“VEV seesaw” [16]) and by the spatial separation of $\Delta_{L}$ from the symmetry-breaking sector. Hence, working in the limit $v_{L}=0$ is well-motivated and does not introduce fine-tuning. Without the loss of generality, the standard type-I seesaw formula, Eq. 18 then gives:
@@
naturally yielding small active neutrino masses without requiring tiny dimensionless couplings. Moreover, in a consistent extra-dimensional framework, the field $\Delta_{R}$ is bulk. Its localization in the extra dimension leads to a universal exponential suppression of all its 44D effective couplings, including its quartic self-coupling. The exponential suppression of the effective quartic coupling $\lambda$ arises from the volume suppression of the 55D interaction. Following Refs. [25], if the $\Delta_{R}$ field is a bulk field with a wavefunction $f_{\Delta_{R}}(y)\propto e^{-M_{\Delta}y}$, the effective 44D coupling is given by:
$\lambda\equiv\lambda_{4D}=\lambda_{5}\int_{0}^{L}dy\,|f_{\Delta_{R}}(y)|^{4}\propto\lambda_{5}\,e^{-4M_{\Delta}L}$ (27)
where $\lambda_{5}$ is the fundamental 55D coupling. This integral is the consequence of integrating the local 55D interaction over the extra dimension. Therefore, a large $M_{\Delta}L$ leads to $\lambda_{4D}\ll 1$, satisfying the condition for type-I strings.
@@
The exponential suppression $e^{-M_{i}L}$ requires $L>1/M_{i}$ to be significant. However, to ensure the validity of the 44D effective field theory (EFT) at the temperature of the $SU(2)_{R}\times U(1)_{B-L}$ phase transition ($T\sim v_{R}$), we must impose that the compactification scale, $M_{c}\sim 1/L$ is significantly higher than $v_{R}$. This avoids Kaluza-Klein mode excitations and higher-dimensional operators disrupting the dynamics of CS formation. Consequently, we require
$v_{R}\ll M_{c}<M_{i}.$ (28)
This hierarchy ensures that the 44D description of the phase transition and the resulting topological defects remains robust. The small Yukawa couplings are generated by the large bulk mass terms $M_{i}$, not an excessively large extra dimension that would compromise the EFT.
The split seesaw mechanism accommodates the full neutrino flavor structure through the introduction of three RH neutrinos with different bulk mass parameters $M_{i}(i=1,2,3)$, as shown in Ref. [25, 2]. Each RH neutrino has a hierarchically exponentially suppressed $e^{-M_{i}L}$ wavefunction that generates the required mass-squared differences $\Delta m_{21}^{2}$ and $|\Delta m_{31}|^{2}$ which accommodates the observed neutrino oscillation, while the structure of the 5D Yukawa matrix $Y_{i\alpha}^{(5D)}$ determines the PMNS mixing angles. As demonstrated in Ref. [2], this framework can successfully reproduce neutrino oscillation data when implemented with flavor structures. While our CS analysis (Sec. III) considers a common breaking scale, $v_{R}$, the presence of multiple RH neutrinos with different flavors $M_{i}$ does not affect the string formation since the string tension is flavor blind. The topological defect formation depends only on the gauge symmetry breaking pattern and vacuum manifold. Although different $M_{i}$ could lead to varied particle emission channels, in our parameter space the dominant energy loss mechanism is gravitational radiation, which remains flavor-independent.
@@
In Fig. 1, we have shown the neutrino mass evolution in the split seesaw mechanism versus the symmetry-breaking scale. The colored shaded region indicates the NANOGrav [4] excluded regime as illustrated in Fig. 3 via frequency bins. The inside colored curves indicate the different values of the exponential suppression factor. The colored stars correspond to the atmospheric scale mass, $M_{\nu}\simeq 0.05$ eV for each exponential suppression factor for the atmospheric scale mass. The colliders constrain the lower value $v_{R}\gtrsim 10^{5}$ GeV [33] presented by the gray shaded regime. It is worth noting that in the split seesaw mechanism, though the smallness of neutrino mass is gained by the exponential suppression of Yukawa couplings as given in eq. 26 but since the breaking scale $v_{R}$ is observationally constrained for a given value of $\beta$ it puts an upper bound on the suppression factor as shown in the Fig. 1. Therefore, the formation of CSs in LRSM combined with neutrino data, constrains the full mass parameter space of the model.
@@
Figure 1: Neutrino mass evolution versus symmetry-breaking scale has been presented. The colored shaded region indicates the NANOGrav [4] excluded regime as shown in Fig. 3. The colored curves represent the different values of the exponential suppression factor. The colored stars correspond to the atmospheric scale mass, $M_{\nu}\simeq 0.05$ eV for each exponential suppression factor. The colliders constrain the lower value $v_{R}\gtrsim 10^{5}$ GeV [33], shown as the gray shaded regime.